OpenStudy (anonymous):

y=x^2+5x+6 is the vertex -5/2 what is the ordered pair ?

OpenStudy (zpupster):

-b/2a that finds and the x coordinate -5/2(1) = -5/2 we know x=-2.5 we need y to get the second half of the ordered pair plug in x y=(-2.5)^2 + 5(-2.5) +6 y= .25 so ( vertex ordered pair is (-5/2, 1/4)

OpenStudy (anonymous):

so the vertex is -5/2 and the ordered pair is 1/4?

OpenStudy (zpupster):

you only found the x of the x,y pair the vertex is an xy pair -b/2a finds the x and what they call the axis of symmetry

OpenStudy (zpupster):

to get the ordered pair plug in the x into the equation. if youlet x be 0 you will get the y intercept letting y be 0 you will get the x intercpt the graph is a parabola also it points upward becasue a is positive hope that helps,

OpenStudy (anonymous):

that helps a little, thank you